# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

### v2

@@ A.4 Proof of regret analysis
Our framework uses EvoDPO, which updates the current reference policy to the previous KL-regularized optimal policy (only for EvoDPO; if it is DPO, $\pi_{ref,t}=\pi_{ref}$). Thus, $\pi_{ref,t}=\pi^{kl}_{t-1}$. Since the third term $D_{KL}(\pi^{kl}_{t}||\pi^{kl}_{t-1})$ is a non-negative term, this term is dropped out for the upper bound.
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\leq\beta\bigl(D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})\bigr)$
For cumulative regret analysis,
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})-\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{1})\dots+\beta D_{KL}(\pi^{*}_{T}\|\pi^{kl}_{T-1})-\beta D_{KL}(\pi^{*}_{T}\|\pi^{kl}_{T})$
Thus, this term is simplified again
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]$
We analyze the second term of the RHS. Since the deterministic policy is $\pi^{*}$ that selects action $y^{*}$, $D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})$ is simplified as below
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})=\sum_{y_{t}}\pi^{*}_{t}(y^{*}_{t}|x_{t})\log\frac{\pi^{*}_{t}(y^{*}_{t}|x_{t})}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}=1\cdot\log\frac{1}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}$
$D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})$ term is also simplified in same way
$D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})=1\cdot\log\frac{1}{\pi^{kl}_{t-1}(y^{*}_{t-1}|x_{t})}$
Then
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})=\log\frac{\pi^{kl}_{t-1}(y^{*}_{t-1}|x_{t})}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}$
We take the definition of each numerator and denominator again
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\leq\log\frac{\pi_{ref,t-1}(y^{*}_{t-1}|x_{t})}{\pi_{ref,t-1}(y^{*}_{t}|x_{t})}+\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}$
From assumption 2, whenever a switch happens,
$\log\pi_{ref,{t-1}}(y^{*}_{t-1}|x_{t})-\log\pi_{ref,t-1}(y^{*}_{t}|x_{t})\leq\log 1-\log\pi_{min}\leq\log\frac{1}{\pi_{min}}$
If a switch has not happened, the $\log$ term is exactly 0. Thus, this $\log$ term follows the cumulative reference mismatch term, which is described in assumption 3; the total number of oracle-optimal action switch bound condition
$\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]\leq B_{T}\cdot\log\frac{1}{\pi_{min}}$
Since the second term $\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}$ is environment drift term and cumulated drift is bounded by $V_{T}$,
$\sum^{T}_{t=1}\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}\leq\frac{\phi_{max}}{\beta}V_{T}$
From the Lemma A.3, we derived the upper bound of $B_{T}$
$B_{T}\leq\frac{2\phi_{max}}{\gamma}V_{T}$
Since the cumulative reference mismatch bias term in total time $T$ is
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]$
$\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\biggl(\frac{2\phi_{max}V_{T}}{\gamma}\cdot\log\frac{1}{\pi_{min}}+\frac{\phi_{max}V_{T}}{\gamma}\biggr)$
Therefore, the cumulative reference mismatch bias term is
$R^{bias}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\mathcal{O}(V_{T})$
@@
Combining the learning error with the reference bias we obtain:
$R_{T}=R^{error}_{T}+R^{bias}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa}+V_{T})\quad w.p\quad 1-\delta.$
Appendix B Additional Information of Evaluation Setup for Section 3
B.1 Detailed Method Used Formulations
To validate the effectiveness of ATLAS, we compare our proposed method against a standard single-agent baseline:
EvoTune (Baseline): This configuration represents the standard self-evolving agent without external supporter agent guidance. The EvoTune generates solutions via standard DPO (mode=vanilla) on the self-generated data. Crucially, EVOTUNE_USE_TEACHER=0 is set, meaning the Exploration Supporter is disabled (no static analysis feedback) and the Strategist is disabled (use_dpo_strategy_teacher=0), forcing the agent to rely solely on the scalar scores from the evaluator.
EvoDPO (Ablation): This isolates the impact of our proposed Evolving DPO (EvoDPO) algorithm (mode=adaptive) without the full agentic support of ATLAS. Unlike EvoTune (Baseline), the reference model $\pi_{\text{ref,t}}$ is not static; it updates to track the evolving policy $\pi_{t}$. However, unlike ATLAS, the Policy Inspector is disabled (or set to always accept), and the Supporters are disabled (EVOTUNE_USE_TEACHER=0), meaning the reference model updates without KL trust-region safeguard and qualitative feedback. This serves as an ablation to test whether the ”moving reference” mechanism alone is sufficient without the Policy Inspector gating.
@@ B.1 Detailed Method Used Formulations
B.2 Detailed Problem Formulations and Evaluation Metrics
B.2.1 Non-Stationary Contextual Bandit
@@
Evaluation. The Research Agent simulates a bandit problem for each candidate policy. The system evaluates performance using the Negative Mean Regret (NMR) over the horizon $H$. This scoring metric aligns with a maximization objective (higher is better):
@@
where $\mathcal{L}_{pde}$, $\mathcal{L}_{ic}$, and $\mathcal{L}_{bc}$ are the mean squared errors (MSE) of the PDE residuals, initial condition loss, and boundary condition loss, respectively. The Research Agent’s task is to distribute the weighting coefficients $\lambda_{pde},\lambda_{ic},\lambda_{bc}$ to balance these competing objectives.
@@
A key challenge in PINN training is the dependency on the Reynolds number $Re\propto 1/\nu$. Low Reynolds numbers (high $\nu$) correspond to diffusion-dominated regimes where solutions are smooth and easy to train. High Reynolds numbers (low $\nu$) correspond to convection-dominated regimes where sharp gradients (shocks) form, creating a “stiff” optimization landscape that often leads to failure modes like the trivial zero solution or spectral bias. To evaluate the research agent’s ability to adapt to these changing physical regimes, we implement a drifting viscosity schedule:
@@
where $\nu_{base}\approx 0.00318$ ($Re\approx 600$), $\alpha=0.9$. This drifts the system between a viscous regime ($\nu\approx 0.006,Re\approx 300$) and a highly turbulent-like shock regime ($\nu\approx 0.0003,Re\approx 6600$). This forces the Research Agent to design a loss strategy that is robust not just to a single static problem, but to the entire transition from smooth flow to shock formation.
@@
B.3 Detailed Role Definition
B.3.1 Non-Stationary Contextual Bandit
Self-Evolving Research Agent.
@@ B.3.1 Non-Stationary Contextual Bandit
The Research Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Research Agent in online learning and contextual bandits. (see the adaptation details in Appendix C.3). For each round, it is provided with the problem context and instructed to implement a Python function policy(context, history, t, ...) that selects an action index (to select arm) $a_{t}\in[0,K-1]$ which will be evaluated by the generated policy and returns the score.
Task-Distributed Multi-LLM Supporters.
@@
The Exploration Supporter (gpt-oss:120b) analyzes the generated policy code. Specifically, the Sliding-Window LinUCB implementation dynamically suggests hyperparameter adjustments. For instance, while the initial window size is set to 50, the supporter may suggest increasing it (e.g./ to 100 or 200) during stable phases to capture longer-term dependencies, or reducing it to adapt more rapidly during high-drift phases. Similarly, it tunes the regularization parameter (e.g., from 1.0 to 0.1) and exploration scalar based on recent performance. In one instance, it flagged a policy with a large fixed window: “The default sliding-window size ($\approx 50$) is too large for a rapidly drifting environment, so the ridge estimator incorporates stale data.” It explicitly patched the code to: if t == 0: policy.window_size = 20 policy.lambda_reg = 0.1 This reduced the estimator’s ”memory” to track the random walk $\theta_{t}$ more effectively.
@@
The Fine-Tuning Strategist (deepseek-r1:32b) monitors training diagnostics during the periodic finetuning updates (every 20 rounds) to adjust DPO hyperparameters (score threshold, $\beta$, and epochs) to prevent the model from overfitting to stale bandit feedback. When it was observed that the student was generating repetitive code (low entropy), it issued a JSON command to set the DPO beta ($\beta$) to $0.1$ and adjust the score_threshold to include a broader range of ”good enough” examples, forcing the model to explore alternative ridge hyperparameters.
@@
The Policy Inspector (llama4:latest) acts as a safety gate for EvoDPO, permitting reference policy updates only when the candidate $\pi^{kl}_{t}$ shows sufficient score improvement ($\Delta_{t}\geq-0.0007$) and satisfies a trust-region KL divergence constraint ($\widehat{\mathrm{KL}}_{k}(\pi^{kl}_{k}||\pi_{\text{ref},k})\leq 0.002$). The logs show it rejecting a candidate update that improved training regret but had a high KL divergence of $0.0045$ on the generated tokens, preventing a ”catastrophic update” where the agent would have dissolved its stable baseline policy in favor of a high-variance experimental one.
B.3.2 Physics-Informed Neural Network (Burgers’ Equation)
Self-Evolving Research Agent.

### v3

@@ B.4 Dynamic regret decomposition
Our framework uses EvoDPO, which updates the current reference policy to the previous KL-regularized optimal policy (only for EvoDPO; if it is DPO, $\pi_{ref,t}=\pi_{ref}$). Thus, $\pi_{ref,t}=\pi^{kl}_{t-1}$. Since $D_{KL}(\pi^{kl}_{t}||\pi^{kl}_{t-1})\geq 0$, dropping it yields the upper bound,
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\leq\beta_{\mathrm{ref}}\bigl(D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})\bigr).$
Summing over $t=1,\ldots,T$ gives
$R_{T}^{\mathrm{bias}}\leq\beta_{\mathrm{ref}}\sum_{t=1}^{T}\left[D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t-1}^{\mathrm{kl}})-D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t}^{\mathrm{kl}})\right].$
Because the oracle policy $\pi_{t}^{*}$ changes over time, this expression does not telescope directly. We therefore add and subtract $D_{\mathrm{KL}}(\pi_{t-1}^{*}\|\pi_{t-1}^{\mathrm{kl}})$ for $t\geq 2$:
$R_{T}^{\mathrm{bias}}\leq\beta_{\mathrm{ref}}D_{\mathrm{KL}}(\pi_{1}^{*}\|\pi_{0}^{\mathrm{kl}})+\beta_{\mathrm{ref}}\sum_{t=2}^{T}\left[D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t-1}^{\mathrm{kl}})-D_{\mathrm{KL}}(\pi_{t-1}^{*}\|\pi_{t-1}^{\mathrm{kl}})\right].$
For deterministic oracle policies,
$D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t-1}^{\mathrm{kl}})=\log\frac{1}{\pi_{t-1}^{\mathrm{kl}}(y_{t}^{*}(x_{t})\mid x_{t})}.$
Therefore,
$D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t-1}^{\mathrm{kl}})-D_{\mathrm{KL}}(\pi_{t-1}^{*}\|\pi_{t-1}^{\mathrm{kl}})=\log\frac{\pi_{t-1}^{\mathrm{kl}}(y_{t-1}^{*}(x_{t})\mid x_{t})}{\pi_{t-1}^{\mathrm{kl}}(y_{t}^{*}(x_{t})\mid x_{t})}.$
If $y_{t}^{*}(x_{t})=y_{t-1}^{*}(x_{t})$, this term is zero. If a switch occurs, the full-support assumption gives
$\log\frac{\pi_{t-1}^{\mathrm{kl}}(y_{t-1}^{*}(x_{t})\mid x_{t})}{\pi_{t-1}^{\mathrm{kl}}(y_{t}^{*}(x_{t})\mid x_{t})}\leq\log\frac{1}{\pi_{\min}}.$
Thus, the cumulative switching contribution is bounded by
$\sum_{t=2}^{T}\left[D_{\mathrm{KL}}(\pi_{t}^{*}\|\pi_{t-1}^{\mathrm{kl}})-D_{\mathrm{KL}}(\pi_{t-1}^{*}\|\pi_{t-1}^{\mathrm{kl}})\right]\leq B_{T}\log\frac{1}{\pi_{\min}}+\frac{\phi_{\max}}{\beta_{\mathrm{ref}}}V_{T}.$
The additional term $\frac{\phi_{\max}}{\beta_{\mathrm{ref}}}V_{T}$ accounts for the utility drift in the exponential-tilted policy between consecutive environments. Combining the switching and drift contributions yields
$R_{T}^{\mathrm{bias}}\leq\beta_{\mathrm{ref}}D_{\mathrm{KL}}(\pi_{1}^{*}\|\pi_{0}^{\mathrm{kl}})+\beta_{\mathrm{ref}}B_{T}\log\frac{1}{\pi_{\min}}+\phi_{\max}V_{T}.$
Using Lemma B.6,
$B_{T}\leq\frac{2\phi_{\max}}{\gamma}V_{T},$
we obtain
$R_{T}^{\mathrm{bias}}\leq\beta_{\mathrm{ref}}D_{\mathrm{KL}}(\pi_{1}^{*}\|\pi_{0}^{\mathrm{kl}})+\frac{2\beta_{\mathrm{ref}}\phi_{\max}}{\gamma}V_{T}\log\frac{1}{\pi_{\min}}+\phi_{\max}V_{T}.$
If the initial KL term is bounded by a constant, then
$R_{T}^{\mathrm{bias}}=\mathcal{O}(1+V_{T}).$
When focusing on the drift-dependent term, this is written as
$R_{T}^{\mathrm{bias}}=\mathcal{O}(V_{T}).$
@@
Combining the learning error with the reference-induced bias gives, with probability at least $1-\delta$,
$R_{T}=R_{T}^{\mathrm{error}}+R_{T}^{\mathrm{bias}}\leq\mathcal{O}\left(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log(T/\delta)}+T^{1-\kappa}+V_{T}+1\right).$
If the initial mismatch term is absorbed into the constant, this is written as
$R_{T}\leq\mathcal{O}\left(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log(T/\delta)}+T^{1-\kappa}+V_{T}\right).$
In particular, the average dynamic regret vanishes when
$V_{T}=o(T^{1-\kappa}),$
so the result is most meaningful in slowly drifting environments.
Appendix C Additional Information of Evaluation Setup for Section 3
C.1 Detailed Method Used Formulations
To validate the effectiveness of ATLAS, we compare our proposed method against a standard single-agent baseline: FunSearch (Baseline): AutoGen (Baseline):
EvoTune (Baseline): This configuration represents the standard Evolving Agent without external supporter agent guidance. The EvoTune generates solutions via standard DPO (mode=vanilla) on the self-generated data. Crucially, EVOTUNE_USE_TEACHER=0 is set, meaning the Exploration Supporter is disabled (no static analysis feedback) and the Strategist is disabled (use_dpo_strategy_teacher=0), forcing the agent to rely solely on the scalar scores from the evaluator.
EvoDPO (Ablation): This isolates the impact of our proposed Evolving DPO (EvoDPO) algorithm (mode=adaptive) without the full agentic support of ATLAS. Unlike EvoTune (Baseline), the reference model $\pi_{\text{ref,t}}$ is not static; it updates to track the Evolving Agent $\pi_{t}$. However, unlike ATLAS, the Policy Inspector is disabled (or set to always accept), and the Supporters are disabled (EVOTUNE_USE_TEACHER=0), meaning the reference model updates without proxy-KL gate safeguard and qualitative feedback. This serves as an ablation to test whether the "moving reference" mechanism alone is sufficient without the Policy Inspector gating.
@@ C.1 Detailed Method Used Formulations
C.2 Detailed Problem Formulations and Evaluation Metrics
C.2.1 Non-Stationary Contextual Bandit
@@
Evaluation. The Evolving Agent simulates a bandit problem for each candidate policy. The system evaluates performance using the Negative Mean Regret (NMR) over the horizon $H$. This scoring metric aligns with a maximization objective (higher is better):
@@
where $\mathcal{L}_{pde}$, $\mathcal{L}_{ic}$, and $\mathcal{L}_{bc}$ are the mean squared errors (MSE) of the PDE residuals, initial condition loss, and boundary condition loss, respectively. The Evolving Agent’s task is to distribute the weighting coefficients $\lambda_{pde},\lambda_{ic},\lambda_{bc}$ to balance these competing objectives.
@@
A key challenge in PINN training is the dependency on the Reynolds number $Re\propto 1/\nu$. Low Reynolds numbers (high $\nu$) correspond to diffusion-dominated regimes where solutions are smooth and easy to train. High Reynolds numbers (low $\nu$) correspond to convection-dominated regimes where sharp gradients (shocks) form, creating a “stiff” optimization landscape that often leads to failure modes like the trivial zero solution or spectral bias. To evaluate the Evolving Agent’s ability to adapt to these changing physical regimes, we implement a drifting viscosity schedule:
@@
where $\nu_{base}\approx 0.00318$ ($Re\approx 600$), $\alpha=0.9$. This drifts the system between a viscous regime ($\nu\approx 0.006,Re\approx 300$) and a highly turbulent-like shock regime ($\nu\approx 0.0003,Re\approx 6600$). This forces the Evolving Agent to design a loss strategy that is robust not just to a single static problem, but to the entire transition from smooth flow to shock formation.
@@
C.2.3 Traveling Salesperson Problem (TSP)
We consider Euclidean TSP instances defined on a complete undirected graph $G=(V,E)$, where $V=\{1,\ldots,n\}$ denotes the node set and each edge $(i,j)\in E$ has distance $d_{ij}$. A candidate solution is represented by a permutation $\pi$ of the nodes. The corresponding tour length is
$L(\pi)=\sum_{i=1}^{n-1}d_{\pi(i),\pi(i+1)}+d_{\pi(n),\pi(1)}.$ (17)
The objective is to minimize $L(\pi)$ subject to the constraint that every node is visited exactly once and the tour returns to the starting node.
The Evolving Agent’s task is to generate a valid tour in a machine-readable JSON format. For each candidate tour, a deterministic verifier checks whether the output forms a valid Hamiltonian cycle and computes the tour length using the provided distance matrix. Invalid tours, including tours with duplicated or omitted nodes, are penalized by the evaluator and are not treated as feasible improvements.
Evaluation. For each TSP instance, we compare the candidate tour length $L_{\mathrm{alg}}$ against an exact-solver reference $L_{\mathrm{opt}}$. We use the negative optimality gap as the task score:
$\mathrm{score}=-\frac{L_{\mathrm{alg}}-L_{\mathrm{opt}}}{L_{\mathrm{opt}}}\times 100.$ (18)
This converts the minimization objective into a maximization score, where values closer to 0 indicate better tours. In the main text, we report the relative improvement from the initial policy to the final policy. We additionally track the valid tour rate and JSON format-compliance rate to ensure that performance improvements are not caused by invalid or unparsable outputs.
C.2.4 Bin Packing Problem (BPP)
We consider one-dimensional Bin Packing instances with item weights $\{w_{i}\}_{i=1}^{n}$, where $w_{i}\in(0,C]$, and identical bins of capacity $C$. A candidate solution partitions the items into bins $\{B_{1},\ldots,B_{k}\}$ such that
$\sum_{i\in B_{j}}w_{i}\leq C,\qquad j=1,\ldots,k.$ (19)
The objective is to minimize the number of bins:
$\min k.$ (20)
The Evolving Agent’s task is to generate a feasible bin assignment in JSON format. For each candidate packing, a deterministic verifier checks that every item is assigned exactly once, no item is duplicated, and no bin exceeds the capacity constraint. Candidate outputs with capacity violations, duplicated items, or unassigned items are penalized by the evaluator.
Evaluation. Let $k_{\mathrm{alg}}$ denote the number of bins used by a candidate solution and let $k_{\mathrm{opt}}$ denote the optimal number of bins computed by an integer-programming solver. We define the score as the negative extra-bin count:
$\mathrm{score}=-\left(k_{\mathrm{alg}}-k_{\mathrm{opt}}\right).$ (21)
This again converts the minimization objective into a maximization score, where larger values indicate fewer extra bins relative to the exact reference. In the main text, we report the relative improvement from the initial policy to the final policy. We additionally report constraint violation rate and packing density as auxiliary diagnostics.
C.3 Detailed Role Definition
C.3.1 Non-Stationary Contextual Bandit
Evolving Agent.
@@ C.3.1 Non-Stationary Contextual Bandit
The Evolving Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Evolving Agent in online learning and contextual bandits. (see the adaptation details in Appendix E.5). For each round, it is provided with the problem context and instructed to implement a Python function policy(context, history, t, ...) that selects an action index (to select arm) $a_{t}\in[0,K-1]$ which will be evaluated by the generated policy and returns the score.
Role-Specialized Multi-LLM Training Supporters.
@@
The Exploration Supporter (gpt-oss:120b) analyzes the generated policy code. Specifically, the Sliding-Window LinUCB implementation dynamically suggests hyperparameter adjustments. For instance, while the initial window size is set to 50, the supporter may suggest increasing it (e.g./ to 100 or 200) during stable phases to capture longer-term dependencies, or reducing it to adapt more rapidly during high-drift phases. Similarly, it tunes the regularization parameter (e.g., from 1.0 to 0.1) and exploration scalar based on recent performance. In one instance, it flagged a policy with a large fixed window: “The default sliding-window size ($\approx 50$) is too large for a rapidly drifting environment, so the ridge estimator incorporates stale data.” It explicitly patched the code to: if t == 0: policy.window_size = 20 policy.lambda_reg = 0.1 This reduced the estimator’s "memory" to track the random walk $\theta_{t}$ more effectively.
@@
The Fine-Tuning Strategist (Qwen 3-32b) monitors training diagnostics during the periodic finetuning updates (every 20 rounds) to adjust DPO hyperparameters (score threshold, $\beta_{DPO}$, and epochs) to prevent the model from overfitting to stale bandit feedback. When it was observed that the student was generating repetitive code (low entropy), it issued a JSON command to set the DPO beta ($\beta_{DPO}$) to $0.1$ and adjust the score_threshold to include a broader range of "good enough" examples, forcing the model to explore alternative ridge hyperparameters.
@@
The Policy Inspector (llama4:latest) acts as a safety gate for EvoDPO, permitting reference policy updates only when the candidate $\pi^{kl}_{t}$ shows sufficient score improvement ($\Delta_{t}\geq-0.0007$) and satisfies a trust-region KL divergence constraint ($\widehat{\mathrm{KL}}_{k}(\pi^{kl}_{k}||\pi_{\text{ref},k})\leq 0.002$). The logs show it rejecting a candidate update that improved training regret but had a high KL divergence of $0.0045$ on the generated tokens, preventing a "catastrophic update" where the agent would have dissolved its stable baseline policy in favor of a high-variance experimental one.
C.3.2 Physics-Informed Neural Network (Burgers’ Equation)
Evolving Agent.